two insect A( -3,2,0) and B(1,1,-3) fly towards a point C(0,2,0) at the same speed.which will arrive first
1
Expert's answer
2020-12-14T07:22:46-0500

Answer

For A insect distance

"r_A=C-A=3i"

So magnitude rA=3m

For B insect distance

"r_B=C-B=-i+j+3k"

So magnitude of distance

rB="\sqrt{11}=3.32m"


So insect A reaches first at point C.